OpenStudy (anonymous):

A 71.0kg fireman slides down a 4.73 meter pole by applying 350. N against the pole with his hands. If he slides from rest, how fast is he moving when he reaches the ground?

OpenStudy (anonymous):

i got like 20 minutes to submit this answer

OpenStudy (anonymous):

i'm drowning

OpenStudy (anonymous):

acceleration due to gravity is 9.8

OpenStudy (anonymous):

NetForce=659.8-350=309.8

OpenStudy (anonymous):

divide by his mass we get net acceleration of 4.36m/s^2

OpenStudy (anonymous):

k....now how do we get that into v = ......m/s?

OpenStudy (anonymous):

or is that just the correct answer?

OpenStudy (anonymous):

vf^2= vi^2+ 2a d

OpenStudy (anonymous):

vf^2= 0 + 2 (4.36)(4.73)

OpenStudy (anonymous):

vf=6.42227

OpenStudy (anonymous):

take the root

OpenStudy (anonymous):

okay...yea i followed that

OpenStudy (anonymous):

I already have

OpenStudy (anonymous):

yea...you did it before i said it lol

OpenStudy (anonymous):

thanks a ton for your help with that

OpenStudy (anonymous):

I hope this is right

OpenStudy (anonymous):

it doesn't matter if it isnt...it's better than i was coming up with

OpenStudy (anonymous):

what were you thinking of?

OpenStudy (anonymous):

well i was doing it right...but didn't even think about gravity

OpenStudy (anonymous):

so i was coming up with weird answers

OpenStudy (anonymous):

and was close to just making an educated guess

OpenStudy (anonymous):

Yeah, always consider gravity unless it is really small. One good thing I notice is they use metric unit in math class now; finally

OpenStudy (anonymous):

it would be a pain if it were in stupid unit